A gas-filled balloon having a volume of 2.90 L at 1.25 atm and 23 degrees Celsius is allowed to rise to the stratosphere (about 30 km above the surface of the Earth), where the temperature and pressure are -44 degrees Celsius and 5.30 x 10-3 atm, respectively. Calculate the final volume of the balloon.
Answer #1

By using combined gas law

P1V1/T1 = P2V2/T2

P1= initial pressure = 1.25 atm ; P2 = final pressure = 5.30*10^-3 atm

V1= initial volume = 2.90 L ; V2 = final volume = ?

T1= initial temperature = 23+273.15= 296.15 K

T2= final temperature = -44+273.15= 229.15 K

V2= P1V1T2/T1P2

= (1.25*2.90*229.15)/(296.15*5.30*10^-3)

529.2 L